If a airplane was travling 1260 miles for 1 hour and 40 minutes what will be the speed?
Answer by jim_thompson5910(35256)

60 min = 1 hour

1 min = 1/60 hour

40 min = 40/60 hour

40 min = 2/3 hour


So 1 hour and 40 minutes = 1 hr + 2/3 hr = 1 2/3 hr = 5/3 hr

-----------------------------------------

d = rt

1260 = r(5/3)

1260*3 = 5r

3780 = 5r

3780/5 = r

756 = r

r = 756


So the speed is 756 miles per hour